OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? 5x + 6y = -5 2x + 3y = -2

OpenStudy (luigi0210):

You could use elimination and multiply the bottom by a -2

OpenStudy (kl0723):

multply the bottom eq by -2 so you can match 5x + 6y = -5 -10x - 6y = 4

OpenStudy (kl0723):

now positive 6y and negative 6y cancel and you are left with one variable

OpenStudy (kl0723):

-5x = -1should be your new eq and now solve for x by isolating it, once that value is found substitute that into one of the eq to obtain the y value